OR2003 Proceedings: Instructions for Authors

1 General comments

The proceedings of OR2003 will be produced by Springer from camera-ready manuscripts. The preparation of manuscripts which are to be reproduced by photo-offset requires special care. We therefore kindly ask you to read very carefully the instructions given below as well as the documentation which comes along with the Springer style files you are asked to use. Contributions submitted in a technically unsuitable form or which do not obey the formatting instructions, will be returned for retyping, or canceled if the proceedings volume cannot otherwise be finished on time. Moreover, note that no changes can be made to the manuscripts after they have been sent to the publisher. You will not receive galley proofs. This means that particular care should be invested to proof-read your manuscript. Please do not forget to apply a spell-checker to your manuscript before sending it away.

Those among you who have submitted contributions to OR proceedings prior to 2001, are asked to take into account that the instructions have changed considerably. It is not any longer acceptable to prepare your manuscript with your favourite program and only take care of the sizes of the typing area and the margins. Please make sure to follow the guidelines given below.

2 Format of submissions

Authors may choose between two alternative procedures to prepare their manuscript, the first of which makes use of LaTeX and the second of Microsoft Word. We recommend you to use LaTeX whenever possible. This will save yourself and the editors a lot of work. If you choose to use LaTeX, please read the instructions for LaTeX contributions. Those using Word are asked to study the instructions for contributions in Word. (You will notice that the manual for Word users is considerably longer than than the manual for LaTeX users, another reason for using LaTeX if possible.)

Please note that manuscripts which are prepared with any system other than LaTeX or Word will be sent back to the authors for retyping/reformating. The same happens to contributions in LaTeX or Word which have not been prepared with the style files/templates provided by Springer.

2.1 Instructions for users of LaTeX2e

Please use the Springer LaTeX2e package for authors/contributors to multi-author books and proceedings. Note that this package works for both English and German contributions. The documentation is available only in English. We suggest that you proceed as follows:

1. Download the package in one of the two ways indicated below (either from the conference web site, or directly from the Springer web site). After downloading you should have the following files

    cl2emult.cls
    cropmark.sty
    math.sty
    sprmindx.sty
    subeqnar.sty
    author.tex
    author.dvi

author.ps (called 2demo.ps at the Springer web site)

    figure.eps
    1readme.pdf

1readme-econ.pdf (only available at the conference web site)

The first five files are the style files needed to prepare your manuscript. The next four files are sample files where author.tex is showing how the style files are to applied. The files author.dvi and author.ps are the corresponding Dvi and PostScript files which result from processing the LaTeX file author.tex. The file figure.eps contains a sample figure file.

The two readme files contain the documentation of the package. The second readme file (file 1readme-econ.pdf) is a shorter readme file specifically designed by Springer for Economics publications (such as the OR proceedings).

2. Read the documentation (files 1readme.pdf and/or 1readme-econ.pdf). There all the files included in the package are explained. Moreover, you will find there instructions on how to use the style files as well as a sample file.

3. Go ahead with preparing your manuscript. It is recommendable to start from the sample file author.tex and adapt this file to your purposes. Please do not change the page layout and other settings. Restrict your own definitions to a strict minimum and do not use separate macro files. Your contribution has to consist of a single LaTeX file, possibly accompanied by separate figure files (in eps format). Make sure that the quality of the figures is good enough for photo-offset reproduction.

If you experience problems with the style files, start with consulting the readme files and the sample file. If this does not help, ask your local LaTeX expert for help. If this fails too, feel free to contact the support staff of Springer (recommended for technical questions) or the editors.

2.1.1 Download method 1 (via the conference web site)

The package for LaTeX2e users is available in compressed form in two files formats, as Winzip archive (recommended for Windows users) and as Unix tar file compressed with gzip. To download one of these archives (recommended method) or all files of the package separately direct your browser to

http://or2003.awi.uni-heidelberg.de/registration/proceedings/fr latexpack.html

2.1.2 Download method 2 (via the Springer web site)

To download the package for users of LaTeX2e directly from the Springer web site direct your browser to the address

   ftp://ftp.springer.de/pub/tex/latex/math/proc/author/

Either download all files separately, or if you wish to download all files at once, download the archive math-proc.zip (Winzip format) from

   ftp://ftp.springer.de/pub/tex/latex/math/proc/author/math-proc.zip

Note that the Springer web site contains only the longer version of the readme file (file 1readme.pdf). If you wish to obtain also the shorter one, specifically designed for Economics publications, download the file 1readme-econ.pdf from the conference web page (see the description above).

2.2 Instructions for users of Word

Please use the Springer Word templates for authors/contributors to multi-author books and proceedings. These templates are available in two versions, for English and for German contributions. The template is guaranteed to work for Winword 7 (Win 95), Word 97, Word 2000 and Word for Macintosh. We suggest that you proceed as follows.

1. Download the support package for Word in one of the two ways indicated below (either from the conference web site, or directly from the Springer web site).

• English version: After downloading the English version you should have the following files

            svcrproc-e.dot
            readme-crproc-e.pdf
            crau-engl-instruct.pdf

The first file is the Word template file needed to prepare your contribution. The second file contains a short explanation on how to install and use the template. A much more detailed documentation is contained in the third file.

• German version: If you have downloaded the German version you should have the following files

            svcrproc.dot
            readme-crproc.pdf
            install-svcrproc.doc
            crau-deutsch-anleit.pdf

The first file is the Word template file needed to prepare your contribution. The second and third files contain a short explanation on how to install and use the template. A much more detailed documentation is contained in the fourth file.

2. Read the documentation very carefully. Start with the files describing the installation of the template. Then install the template. Do not forget to adjust the margins with the command change margins as described in the documentation.

Before starting to type your manuscript, please study also the guide for preparing Springer Word documents available as file crau-engl-instruct.pdf (English version) or crau-deutsch-anleit.pdf (German version). Please note that the template and the documentation do not have been set up specifically for the OR proceedings series, but for general multi-authored books and proceedings in mathematics and economics. Some comments and explanations in the documentation do not apply to the case of the OR proceedings - it should not be difficult to identify those (they usually refer to multi-authored books). If you are in doubt, feel free to contact the editors.

If you do not read the guidelines, you can almost be sure that your contribution will not fulfill the rules and will have to be sent back to you. If you do not like to study a long documentation, you might consider using LaTeX as an alternative (or possibly, ask another person to type your manuscript with the help of LaTeX).

3. Go ahead with preparing your manuscript.

Please do not change the page layout and other settings. All instructions given in the documentation have to be obeyed. Your contribution has to consist of a single file; figures should be integrated into your file. All figures must be available electronically. Make sure that the quality of the figures is good enough for photooffset reproduction.

If you experience problems with using the template, start with consulting the documentation files. If this does not help, ask your local Word expert for help. If this fails too, feel free to contact the support staff of Springer (recommended for technical questions) or the editors.

2.2.1 Download method 1 (via the conference web site)

The package for Word users is available in compressed form (Winzip archive) in two language versions (German and English).

To download one of these archives (recommended method) or all files of the package separately direct your browser to

http://or2003.awi.uni-heidelberg.de/registration/proceedings/fr wordpack.html

2.2.2 Download method 2 (via the Springer web site)

• The English version of the package is available from the Springer web site at the address

http://www.springer.de/economics/authors/index.html

Download the files

http://www.springer.de/author/crau-engl-instruct.pdf

and

http://www.springer.de/author/sv-crproc-e.zip (archive containing 2 files)

• The German version of the package is available from the Springer web site at the address

http://www.springer.de/economics-de/authors/index.html

Download the files

http://www.springer.de/author-de/crau-deutsch-anleit.pdf

and

http://www.springer.de/author-de/sv-crproc.zip (archive containing 3 files)